Jolina and Judy Ann can encode a term paper in 3 3/5 hours. Jolina can do it by herself in 6 hours. How long would it take Judy Ann to encode the term paper by herself?
Answer by checkley77(12844)

6X/(6+X)=3.6
6X=3.6(6+X)
6X=21.6+3.6X
6X-3.6X=21.6
2.4X=21.6
X=21.6/2.4
X=9 HOURS FOR JUDY TO ENCODE BY HERSELF.
PROOF:
6*9/(6+9)=3.6
54/15=3.6
3.6=3.6